1 Here is some gray text Here is a lightgray line: And a dark gray blob:

Left Down Right Big Flip

This runs diagonally down and to the right This runs up to the right

\begin{aligned}& A \\ B & C \end{aligned}_{f}_{g}_{h}

1 GOAT

GOAT

Doom

Zoom Polar Coor Polar Coor H M r l Doom